OpenStudy (anonymous):

Will medal and Fan If the quadratic formula is used to solve 2x 2 = 8x - 3, what are the solutions?

TheSmartOne (thesmartone):

\[2x^2-8x+3\]

OpenStudy (anonymous):

Do you need the answer choices?

TheSmartOne (thesmartone):

Sure.

TheSmartOne (thesmartone):

You could plug in the answer choices into the equation to see which ones work.

TheSmartOne (thesmartone):

\[8\pm \sqrt{64-4(3)(2)}\div4\]

OpenStudy (anonymous):

file:///Users/nolanhairr/Desktop/Screen%20Shot%202014-09-24%20at%206.03.11%20PM.png

TheSmartOne (thesmartone):

\[8\pm \sqrt{40}\div4\] \[\frac{ 8\pm2\sqrt{10} }{ 4 }\] \[\frac{ 4\pm \sqrt{10} }{ 2 }\]